OpenStudy (anonymous):

Test soon! Calc optiminaztion problem. Please help! A French aquarium is to be built with a rectangular slate bottom twice as long as it is wide, and with four rectangular glass sides. There is no top. Slate for the bottom costs $2 per square meter while the glass used cost $4 per square meter. Find the dimensions of the cheapest tank having a volume of 18 cubic meters.

OpenStudy (welshfella):

let x be width of the base then 2x is the length let h be depth so V = 2x*x*h = 2x^2h 18 = 2x^2 h h = 18/2x^2 = 9 / x^2

OpenStudy (welshfella):

now find an expression for the cost of the material are of the base = 2x^2 so cost will be 2*2x^2 = 4x^2 area of the 4 sides = 2hx + 4hx = 6hx so cost = 4 * 6hx = 24hx

OpenStudy (welshfella):

now you can plug in h = 9 / x^2 into the formula for the cost differentiate with respect to x , equate to zero and find value of x for minimum cost

OpenStudy (welshfella):

Oh they've gone!